OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2. y = 4x - 1 y = 4x - 3 y = 1 fourthx - 3 y = 1 fourthx - 1

OpenStudy (bradely):

y=(1/4)x+2 Slope m=1/4 Slopes of the parallel lines are equal m=1/4 and (-4,-2) y=(1/4)x+c -2=(1/4)(-4)+c -2=-1+c c=-1 y=(1/4)x-1 Option 4 is correct Source:www.mathskey.com

OpenStudy (anonymous):

thank You, can you help me with another one.

OpenStudy (bradely):

please post source site

OpenStudy (anonymous):

It is homework

OpenStudy (bradely):

yes

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (bradely):

Slope of the line is -1 Perpendicular line slope ,m =-1/-1=1 y=x+c Point (-4,-2) -2=-4+c c=2 y=x+2 option 4 is correct http://www.mathskey.com/question2answer/ask